OpenStudy (anonymous):

Two students, Jennifer and Jamal, factored the trinomial 6x2 − 3x − 9. Jamal factored it as 3(2x − 3)(x + 1) and Jennifer factored it as (2x − 3)(3x + 3). Indicate which student factored the trinomial completely and which student did not, and explain why

pooja195 (pooja195):

In order to figure out who did it wrong we need to solve it outselves our a=6 b=-3 c=-9 values will help in this case :) Remeber the abc values come from this formula \[\huge~\rm~ax^2+bx+c \]  \[\huge~\rm~a*c=6\times-9\] \[\huge~\rm~6 \times -9=? \]